Operations Report – April 29th

Crew 211 Operations Report 29-APR-2019

SOL: 8

Name of person filing report: Julien Amalaberque

Non-nominal systems: nothing to report

Notes on non-nominal systems: nothing to report

Generator (hours run last night): 13:00

On (last night): 6:30 pm

Off (this morning): 7:30 am

Solar— SOC 71 % at 14:30 pm

Diesel Reading – 25 %

Propane Reading – 49 %

Ethanol Free Gasoline – 4 gallons

Water (auxillary tank) – 0 gallons

Water (static tank) – ~290 gallons

Auxillary to Static tank transfer– no

Gallons transferred: No

Water in GreenHab – 80 gallons

Water (loft) – Static to Loft Pump used – yes

Water Meter: 15 gallons

Toilet tank emptied: yes

Deimos rover used: no (out of order)

Hours:

Beginning charge:

Ending charge:

Currently charging:

Sojourner rover used: ASSIGNED TO DIRECTOR, no

Hours:

Beginning charge:

Ending charge:

Currently charging:

Spirit rover used: yes

Hours: 90.2h

Beginning charge: 100

Ending charge: 55

Currently charging: yes

Opportunity rover used: no (out of order)

Hours:

Beginning charge:

Ending charge:

Currently charging:

Curiosity rover used: yes

Hours: 92.2h

Beginning charge: 100

Ending charge: 72

Currently charging: yes

Notes on rovers: Nothing to report

Reason for use: EVA with PBS

ATV Fuel Used: 0 Gals

# Hours the ATVs were Used today: 0

Notes on ATVs: nothing to report

HabCar used and why, where? No

CrewCar used and why, where? No

General notes and comments: Nothing to report

Summary of internet: Nothing to report

Summary of suits and radios: Nothing to report

Summary of Hab operations: Nothing to report

Summary of GreenHab operations: Some basil and parseleaf was harvested

Summary of ScienceDome operations: Nothing to report

Summary of RAM operations: Nothing to report

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support:

* Some additional water for the GreenHab would be useful.

Sol Summary – April 29th

Crew 211 Sol 8 Summary Report 29-APR-2019

Sol: 8

Summary Title: Mars on TV

Author’s name: Carl-Henrik Dahlqvist

Mission Status: Nominal

Sol Activity Summary: After a quick breakfast, the commander welcomed the PBS team in spacesuit outside of the Hab. We then made a tour of the MDRS campus with the TV crew, each member of the UCL to Mars team presenting part of the infrastructures and discussing his/her experiment. Half of the crew left then for an EVA with Liz Adeola to Cow Dung Road between the Robert’s Rock Garden and Zurbin’s head. The EVA team got back at noon and we all had a nice lunch together. The crew commander and the crew astronomer were then interviewed for about one hour. All the crew members went then back to work once the TV crew left the campus.

Look Ahead Plan: Depending on the weather, we plan to start the day with an EVA to Kissing Camel. We will then work on our respective experiments for the rest of the day preparing the first experiment related EVA for the commander and vice-commander which should be organized on Wednesday.

Anomalies in work: No issues.

Weather: Showers, mostly cloudy and windy

Crew Physical Status: Good

EVA: Half of the crew explored the west part of Cow Dung Road with the PBS crew. Liz Adeola was part of the EVA team, as we her a spacesuit for the occasion. The EVA lasted for about an hour. The PBS team did a few drone footages and interviewed the EVA team members.

Reports to be filed: Journalist Report, GreenHab Report, Operations Report, Science Report, EVA report, EVA Request, Daily Photo Report

Support Requested: None

GreenHeb Report – April 28th

Crew 211 GreenHab Report – 28-APR-2019

GreenHab Officer: Chloé Peduzzi

Environmental control: Ambient with door open + fan turned up to 2

Shade cloth: 80%

Average temperatures:

Low: 16°C

High: 45°C

Hours of supplemental light: 0

Daily water usage for crops: 4 gallons

Water in Blue Tank: 85 gallons

Time(s) of watering for crops: 11:30

Changes to crops: All plants look healthy. Nevertheless the tomato plant of the GreenHab seems a little stressed (curved leaf), I will water it accordingly over the next couple of days. We have harvested the lettuce (from plants we have brought) for the dinner.

Narrative: Hot day. I have cleaned the floor of the GreenHab to make sure it stays a pleasant place in which to work. Moreover I have regrogrammed the fan, as you asked me, so it will be “on” from 10:00 to 17:30. Is that okay ?

Harvest: 27.5g lettuce (from plants we have brought)

Support/supplies needed: I have regrogrammed the fan, as you asked me, so it will be on from 10:00 to 17:30. Is that okay ?

EVA Report – April 28th

Crew 211 EVA Report 28-04-2019

EVA #7

Author: Simon Collignon

Purpose of EVA: Exploring the geology of White Rock Canyon
Start time: 09:00 am
End time: 10:45 am

Narrative: The aim of this EVA was to explore the White Rock Canyon and gather some soil samples which were successfully retrieved. After the scientific part of this EVA, we’ve looked after ancient river beds which could be a good evidence of liquid water on Mars.
Destination: White Rock Canyon

Coordinates Latitude: 4248000, longitude: 52000 (UTM NAD27 CONUS)

EVA Participants: Simon Collignon, Julien Amalaberque, Nathan Pechon

Mode of travel: Driving and hiking

Vehicles used: Spirit and Curiosity

Greenhab Report – April 27th

Crew 211 GreenHab Report – 27-APR-2019

GreenHab Officer: Chloé Peduzzi

Environmental control: Ambient with door open + fan turned up to 3

Shade cloth: 80%

Average temperatures:

Low: 14°C
High: 46°C

Hours of supplemental light: 0

Daily water usage for crops: 4 gallons

Water in Blue Tank: 90 gallons

Time(s) of watering for crops: 11:00

Changes to crops: All plants look healthy. Green chard, cucumber (03/20) and romaine lettuce grow very well. Tomatoes (from plants that we brought), carrots and radishes are ready to be harvested. Onion plants are emerging. More flowers are visible on the tomato plant of the GreenHab.

Narrative: Still hot today. No harvest. I’m still watering each morning the bean and tomato seeds of my experiment placed in the GreenHab.

Harvest: None

Support/supplies needed: None